The Efficient Software Store had been selling a spreadsheet program at a rate of 100 per month and a graphics program at the rate of 50 per month. In September 2012, Efficient's supplier lowered the price for the spreadsheet program, and Efficient passed on the savings to customers by lowering its retail price from $400 to $350.The store manager then noticed that not only had sales of the spreadsheet program risen to 120, but also the sales of the graphics program increased to 56 per month. Explain what has happened. Use both arc price elasticity and arc cross-elasticity measures in your answer.


1
Expert's answer
2021-10-03T14:16:30-0400

price elasticity of demand

"=\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}\div
\frac{P_2-P_1}{\frac{P_2+P_1}{2}}"

where Q1 and Q2 are quantity demanded at price P1 and P2 respectively.

Q1=100, Q2=120, P1=400, P2=350

"=\frac{120-100}{\frac{120+100}{2}}\div
\frac{350-400}{\frac{350+400}{2}}\\
=\frac{20}{110}\div\frac{-50}{375}\\=-1.36"

spread sheet is price elastic so decrease in price will rise the sales.


cross elasticity of demand

"=\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}\div
\frac{P_2-P_1}{\frac{P_2+P_1}{2}}"

where Q1 and Q2 are quantity demanded at price P1 and P2 respectively.Q1 and Q2 are quantity demanded for related product

Q1=50, Q2=56, P1=400, P2=350

"=\frac{56-50}{\frac{56+50}{2}}\div
\frac{350-400}{\frac{350+400}{2}}\\
=\frac{6}{53}\div\frac{-50}{375}\\=-0.84"

Decrease in price of graphics will increase demand for graphics because they are complementary goods.